OpenStudy (anonymous):

8. Sand pours from a chute and forms a conical pile whose height is always 2 times the base radius. If the base radius of the pile increases at a rate of 7 feet/hour, find the rate of change of the volume of the conic pile, when the base diameter of the pile is 10 feet.

OpenStudy (anonymous):

wait for a while. i'm solving it.

OpenStudy (anonymous):

\[V=1/3\pi r ^{2}h\] \[dv/dt=1/3\pi rdr/dt (2)h\] \[dv/dt=2/3\pi r dr/dt h\] r=5,h=10, dr/dt=7ft/hr. \[dV/dt=(700\pi)/3\]

OpenStudy (anonymous):

the unit is feet per hour.

OpenStudy (anonymous):

i hope this can help.